OpenStudy (anonymous):

(-6,-1) (-3,2) (-1,4) (2,7) are considered a exponential function right?

hero (hero):

?

OpenStudy (anonymous):

I'm dealing with Linear, Quadratic and Exponential Models.

hero (hero):

4 points alone can't be considered an exponential function though.

hero (hero):

You should clarify your question a bit more.

OpenStudy (anonymous):

here is the formula for an exponential function: f(x)=a(b)^x or f(x)=p(1+r)^x

OpenStudy (anonymous):

Okay, it's kind of hard to explain but, simply put, could you tell me if the x-values and the y-values share the same difference and or ratios?

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

are you asking what the rations are??

hero (hero):

Open Study works pretty simply. All you have to do is post the question you are actually working on. Avoid posting partial questions or areas where you are stuck. Just post the ORIGINAL original ACTUAL question you are working on. Then from there, we can discuss what you are working on, where you are stuck, what you need help with etc.

OpenStudy (anonymous):

x = -6, -3, -1, 2 y = -1, 2, 4, 7 I'm looking at it right now and the x values don't look like they share any same difference. Same for the y values.

OpenStudy (anonymous):

I greatly apologize, but this is the question.

OpenStudy (anonymous):

From -6 to -3, what do you do? Add 3, so that's the rule -- to add 3. But if you add 3 to -3 it's 0, not -1.

OpenStudy (anonymous):

If you can't answer then I'm sorry for wasting your time :/ I'm just confused.

hero (hero):

What is it that you have to do ultimately? Create an equation based on these points I'm guessing.

OpenStudy (anonymous):

Maybe this will help. I'm trying to find out if those points are one of the following functions. Linear Function: If the y values of a set of data have a common difference, then a linear model would best fit the data. Quadratic Function: If the y values of the data have a constant set of second differences, the data should be represented by a quadratic model. Exponential Function: If the y values have a common ratio instead of a common difference, the data should be represented by an exponential model

hero (hero):

Use the slope formula to determine if there is a common difference between the points

hero (hero):

\[m = \frac{y_2 - y_1}{x_2 - x_1}\]

hero (hero):

\[m = \frac{2 -(-1)}{-3 -(-6)} = \frac{3}{3} = 1\]

hero (hero):

\[m = \frac{4 - 2}{-1 -(-3)} = \frac{2}{2} = 1\]

hero (hero):

So yes, looks like you're dealing with points that represent an exponential function.

hero (hero):

But make sure you check the rest of the points to be 100% sure

OpenStudy (anonymous):

Okay, thank you so much. I'm over here scratching my head wondering what to do.

hero (hero):

Basically, if you've tried looking for a common difference, and haven't found one, the next thing to do is to use the slope formula to look for a common ratio. In this case, the common ratio is 1

hero (hero):

So if we describe the process step wise, it would be this: 1. Determine if the function has a common difference. If so, then the function is linear. If not, proceed to step 2: 2. Determine if the function has a 2nd set of common differences. If so, the function is quadratic If not proceed to step 3: 3. Determine if the function has a common ratio by using the slope formula. If so, then function is exponential.

hero (hero):

If you use the slope formula when comparing each set of points and end up with the same number as a result, then the function has a common ratio.

hero (hero):

For this one, even though "1" is an integer, it can still be written in ratio form.

hero (hero):

I apologize if any of what I am saying is confusing.

hero (hero):

All I can say is 3/3 is a ratio 2/2 is a ratio 3/3 = 2/2 = 1

hero (hero):

I don't know exactly what it is you are confused on.

OpenStudy (anonymous):

It's okay, I understand now.